if a person takes a 5 problem true or false test, what is probability that he scores 100 if he guesses on every problem? (Express your answer as a reduced fraction)
Answer by stanbon(75887)

if a person takes a 5 problem true or false test, what is probability that he scores 100 if he guesses on every problem?
--------
Binomial Problem with n = 5 and p(correct) = 1/2
----
P(x = 5) = (1/2)^5 = 1/32 = 0.03125
-----------------------------------
Cheers,
Stan H.
================